# Consumption Smoothing in Metropolis: Evidence from Working-class Households in Prewar Tokyo

### v12

@@ 4.3 Response to Adverse Health Shocks
The foregoing analyses used the within variations in the earnings to estimate income elasticity, meaning that the identification depends both on the adverse and favorable idiosyncratic shocks. While it is useful to illustrate the overall responses of the households to the shocks, it does not reveal the pathways behind the risk-coping strategies. Next, I focus on temporary illness of household heads that could cause temporary income losses.
@@
Notes: This figure shows the time-series plots of the monthly death counts from influenza (ryūkōsei kanbō) in Tokyo prefecture between January 1918 and December 1920. The death statistics used in this figure were collected based on the comprehensive registration system.
Source: Created by the author using the 1918, 1919, and 1920 editions of the Statistics of Causes of Deaths of the Empire Japan (Statistics Bureau of the Cabinet 1921; 1922; 1923).
@@
I began my analysis by testing whether the illness of the head impacted the earnings. To do so, I decided the exact timing and duration of illnesses of the family members using the Kazoku idō hyō (Section 3.4). Figure 5 shows the number of monthly deaths from influenza (ryūkōsei kanbō) in Tokyo between 1918 and 1920. There appears to have been a moderate spike in early 1919, which is partially captured in the THBS period. In fact, a few THBS budget books indicate that the family members contracted influenza during this period. Five households experienced the illness during the sample period, and illness affecting the heads was concentrated between February and May 1919, when influenza exposure was most prevalent. [^63] Among four cells with health shocks on the heads, two are observed in February, during the peak of the epidemics (Figure 5), and the others are in March and May. In addition, the average duration for the heads is $5.8$ days (Std. Dev. $=3.6$), which is consistent with the fact that Type A influenza caused fever for less than six days (Uyeki et al. 2022). Therefore, it is plausible that influenza was responsible for illnesses during this period, providing a quasi-experimental setting for analyzing to what extent the adverse idiosyncratic shock of illness influences the households’ consumption strategies.
@@
Earnings in a semi-month are based on the work done during the previous semi-month. This means that the head’s earnings were influenced by the number of illness days in the previous semi-month cell. Table 6 presents the results from the regression of the head’s earnings on the number of illness days in the previous semi-monthly cell. While the instantaneous health shock has little impact on the head’s earnings (Column 1), the lagged health shock has a statistically significant effect on their earnings (Column 2). This result is unchanged if I include both variables in the same regression in Column 3, supporting the evidence that the health shocks in the previous cell reduced the earnings of the head. The estimate in Column (2) suggests that a one-day increase in illness decreases the head’s semi-monthly income by approximately $2.4$ yen. Notably, this is close to the average daily wage rate of skilled factory workers (Online Appendix B.6). This reflects the reductions in their salary due to absence: Miyoshi (1989, pp. 383; 388) highlighted that income was reduced due to absence when the head was infected with a disease. Finally, Column (4) adds the health shocks of the other family members into the specification of Column (3). The estimated coefficients on these placebo variables are statistically insignificant, supporting the robustness of the result in Column (3). These results confirm that the head who suffered illness lost earnings. [^64] Given that the average number of illness days was approximately $2.0$, the temporary health shock losses were more than five yen in that month, accounting for approximately $10$% of the average monthly income. Although this is a rough estimate and the losses depend on the severity of the shocks, my result shows that the household income responded to the adverse health shock of illness among the factory workers’ households.
@@
Notes: This table shows the results from the regressions of the head’s earnings on the health shock variables using the semi-monthly panel dataset. A household with two semi-month observations is excluded from the sample in all the regressions because the lagged health shock variables are included. The family size variable, household fixed effects, and time-fixed effects are included in all the regressions. “Health shocks on the head” refers to the number of days that the head suffered illness. “Health shocks on the others” refers to the number of days that family members, excluding the head, were ill. “[level]” indicates the variable of the head’s earnings in the same semi-month cell. “[lag]” indicates the lagged variable (i.e., the shocks measured in the previous semi-month cell). Although the correlation coefficient between the level and lag of the health shock variables for the other family members is substantially high (Pearson’s coefficient $=0.99$), including either of the two variables does not change the finding on the negligible impacts of the health shocks on the other family members.
Following the evidence, I characterize Equation 1 as a structural consumption-smoothing equation by using the first-stage reduced form equation as follows:
@@
where $h_{i,t-1}$ is the number of days of the head’s illness in the previous semi-month cell, $\varpi_{i}$ is the household fixed effect, $\tau_{t}$ is the semi-month fixed effect, and $e_{i,t}$ is a random error term. For the risk-sharing mechanisms analysis under equation 2, the log-transformed disposable income in equation 3 ($y_{i,t}$) is replaced with the head’s earnings ($\tilde{y}_{i,t}$). The health shock variable is plausibly exogenous because the illnesses were from the influenza epidemics, as explained above. [^65] As noted in Section 2.3, the factories did not offer compensation for sick leave because the Factory Act did not specify benefits for workers’ illness (Study Group of the Factory Act 1916). Even in the largest factory in Tsukishima, the Ishikawajima Shipyard, any descriptions of the influenza epidemics are documented in the company’s history (Arai 1930). In fact, Miyoshi (1989) documented the issue that the head’s disease directly caused income loss due to the absence at that time. In addition, there is a correlation between the health shock variable and the head’s earnings at a statistically valid level (Table 6). Both support the evidence that the exogeneity of the instrument and relevance condition hold under this just-identified system.
@@
Column 1 shows the results of equation 1, in which the log-transformed disposable income is replaced with the head’s illness (i.e., the number of days with illness in the previous semi-month cell). The dependent variables are the 11 measures of log-transformed consumption, and all the regressions include the family size control, household-fixed effects, and time-fixed effects. The estimated coefficient on the health shock variable is listed in the sub-column named “Coef.”.
Column 2 shows the results of equation 1, in which the log-transformed disposable income is regarded as the endogenous variable. The head’s illness is an instrumental variable (i.e., excluded exogenous regressor) for the first-stage reduced-form equation (3). The estimated coefficient of the disposable income is listed in the sub-column named “Coef.”. The first-stage $F$-statistics on the health shock variable are listed in the final column.
@@
I first consider a reduced-form specification using the head’s illness as the exposure variable in equation 1 following Cochrane (1991). Column (1) in Table 7 shows the results. Overall, the coefficients on the exposure variable are negative, supporting evidence that adverse health shocks on the heads reduce consumption levels. Next, I explicitly model the structure underlying the relationship between the health shocks and earnings using the reduced-form equation of (3). If the negative impacts of the health shocks on consumption were derived from the income losses, the results from the structural estimation approach correspond to those in Column (1). The estimates from the structural equation are listed in Column (2) of Table 7. As shown, the results are congruent with the reduced-form results in Column (1). Importantly, the estimates are systematically larger than those reported in Table 4. The estimates for total and food consumption categories are $0.543$ and $0.355$, respectively, whereas the baseline estimates are $0.358$ and $0.136$. While this result partly reflects the bias reduction under the instrumental variable estimation, it is also noteworthy that this system is designed to reduce the bias by specifying the channel from adverse (instead of favorable) income shocks. For example, expenses on the allowance for children could be easily cut during unfavorable times but may not respond to favorable shocks. Similarly, the expenses for furniture are likely to be cut, but the households might not buy those frequently, meaning that there are no sufficient cases for reacting to rare health shock events. [^66] In other words, furniture may have only been purchased when the households had extra income. This may partly be reflected in the smaller value of the first-stage $F$-statistics in the furniture subcategory. [^67] Note that the logarithmic transformation leaves the cells with expenses on the furniture. Thus, the reduced-form equation captures the relationship between the health shocks and earnings in the cells with any expenses on furniture. While the first-stage $F$-statitics is $35.7$ in the linear-linear model, the second-stage result is robust against the functional form assumption. This may reflect the asymptotic theoretical advantage in the panel data model under nearly weak instruments case (Staiger and Stock 1997; Cai et al. 2012; Baltagi et al. 2012). The same argument is applicable to the housing category. A similar interpretation can be applied to the transportation category because family travel was not a frequent event and thus, unlikely to coincide with the timing of illness. By contrast, the entertainment expenses were still likely to be cut in case of illness, which might reflect the curtailment of tobacco, theater-going, and the allowance for heads. This may be another explanation for the difference in the results from my baseline results in Table 4.
@@
While the estimates for the clothes and gifts categories now exceed one, the null of equality to the population mean of one is not statistically rejected for both cases. Although the large estimates may be partly due to the operation under the instrument based on rare health shock events, the findings are materially like my baseline results—both categories are sensitive to the shocks as are those of the dispensable categories.
@@
Adverse Health Shock as the Idiosyncratic Shock
@@
Head’s illness Obs. Head’s earning Obs.
@@
Panel A: Savings, insurance, borrowing, and gifts (yen)
@@
Other member’s earnings 0.001 [0.081] 254 -0.001 [0.034] 254
@@
Notes:
Column 1 shows the results of equation 2, in which the head’s earning is replaced with the head’s illness (i.e., the number of days with illness in the previous semi-month cell). The family size control, household fixed effects, and time-fixed effects are included in all the regressions. The estimated coefficients on the head’s illness are listed in the sub-column named “Coef.”.
Column 2 shows the results of equation 2, in which the head’s earning is treated as the endogenous variable. The head’s illness is an instrumental variable (i.e., excluded exogenous regressor) in the first-stage reduced-form equation. The first-stage result is identical to the result listed in Column (2) of Table 6 with the first-stage $F$-statistics of $13.6$. The estimated coefficients on the head’s earnings are listed in the sub-column named “Coef.”.
Panel A: Each net income variable is defined as the difference between income and expenses. “Net savings” refers to the difference between the amount withdrawn and the total deposits. “Net insurance” refers to the amount received from insurance minus the expenses paid for insurance. “Net borrowing” refers to the amount of borrowing minus the debt payments. “Net credit purchase” refers to the amount of credit purchases minus credit redeemed. “Net gifts” refers to the total received amounts, including both pecuniary and non-pecuniary gifts, minus the payments for gifts.
Panel B: “Other member’s earnings” refers to the total income earned by all family members except for the head of the household. “Sales of miscellaneous assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles. All dependent variables are in yen.
@@
I tested the risk-coping strategies for mitigating adverse health shocks in Table 8. Column 1 lists the reduced-form results in the same manner as Table 5. The estimated coefficients on the net income variables are positive in most cases, indicating the households’ compensating strategies against the shocks. The estimate for the credit purchase category is weakly statistically significant, with the largest magnitude. It suggests that the one-day increase in the head’s illness increased the net credit purchases by $1.29$ yen, which accounts for approximately half of their average daily wages ($1.29/2.56$). Column 2 presents the estimates from the structural equation using the head’s earnings as the endogenous variable. A similar but clearer result is obtained—the estimate for the credit purchase category is approximately four times greater than that reported in Column (1) of Table 5. It indicates that a one yen decrease in the head’s earnings increases the net income from credit purchases by $0.55$ yen. This magnitude is consistent with the direct effect of the health shocks on the net credit purchase estimated in Column (1) of Table 8.686868This means that $2.56\times 0.55=1.4$ yen, which is close to $1.3$ yen in Column (1). The estimates for the savings and gifts categories are still negative but now statistically insignificant, supporting the evidence that credit purchase is the primary risk-mitigation strategy for dealing with health shocks. Since withdrawing and receiving gifts to use the money at the retailers takes more time (steps), the households likely prefer to use credit purchases.
Overall, the event-study analysis using adverse health shocks confirms the baseline findings and provides a sharper illustration of consumption-smoothing strategies in the case of adverse idiosyncratic shock events. My results show that the unique available response to idiosyncratic income shocks among the THBS households was to use credit purchases. The available compensation via credit purchases accounts for approximately half of their temporary income losses, which partially contributes to smoothing their food consumption in the short term.
5 Credit System in Local Economy
@@
The foregoing results enlighten the roles of credit institutions among skilled workers’ households during industrialization. As the credit for purchases was provided by local retailers, this suggests the importance of the relationships between the workers and retailers. To the best of my knowledge, however, there is little historical research on credit purchases in prewar Japan. Hence, the details of this credit institution are not yet clear. This may mainly be because the information on credit purchases is personal, as in the case of state-contingent loans in the rural credit market (Udry 1994) and thus, not well documented. Accordingly, I analyze the institution using both the THBS dataset and available official documents.
@@ 5.1 Availability and Premiums: Evidence from the THBS Dataset
Notes: Figure 6a, and 6c illustrate the time-series plots of the average daily and semi-monthly credit purchase and payment, respectively. The blue and red dashed lines indicate the average credit purchases and payments, respectively. Figure 6b and 6d illustrate the time-series plots of the average daily and semi-monthly credit purchase minus payment. The red solid line highlights the zero value. Figures 6a and 6b show the daily data between January 12 and November 11, 1919. Figures 6c and 6d show the semi-monthly series calculated using the daily data between January and November 11, 1919. All figures are based on the 29 THBS households that used any credit purchases.
Source: Created by the author using the THBS sample.
I begin this descriptive analysis by examining the credit purchases included in the THBS dataset. Most THBS households ($29/33$) used any credit purchases during their sample period, confirming that credit purchase was a common practice among factory workers. Figure 6a illustrates the average daily credit purchases and payments. Generally, there is a peak per semi-month in payments, which reflects the practice of making the payments for the credit purchase on or soon after their paydays. The purchase series shows a similar tendency but with lower spikes and larger variance. This indicates that the households first purchased goods for a certain amount on or soon after their paydays, and then used purchased perishable foods such as fish, fruit, and vegetables when they needed them. Figure 6b illustrating the net purchase confirms this behavior: the households used the credit on a daily basis and redeemed it near paydays. Figure 6c shows that the payment slightly exceeded purchase, suggesting that many households could approximately repay their loans. Although the difference between payment and purchase is relatively obvious after August 1919, Figure 6d indicates that this difference is not statistically significant in all the semi-months. [^69] Strictly, payment can include accrued liabilities that should have been paid prior to the previous period. It is, therefore, reasonable that the average payment is greater than the average purchase in most semi-month cells. Note again that the wider variances in the latter sample period are based on the smaller available number of households (Online Appendix Figure B.4).
I then analyzed whether the volatility of earnings restricted credit purchases. Since credit of worker was unobservable, I used the volatility of earnings as a proxy of credit risk in the following empirical specification:
@@
where $I(\cdot)$ is an indicator function, $\overline{CP}$ is the (unit-)average credit purchases, $CV$ is the coefficient of variation for head’s earnings, $\bar{\mathbf{x}}$ is a vector of the (unit-)average family composition variables, and $v$ is a random error term. All variables are calculated using the semi-monthly series. [^70] For example, the coefficient of variation for household $i$ is defined as $CV_{i}=\{\sum_{t}^{T_{i}}(\tilde{y}_{i,t}-\bar{\tilde{y}}_{i,.})/(T_{i}-1)\}^{\frac{1}{2}}/\bar{\tilde{y}}_{i,.}$, where $T_{i}$ indicates the number of time-series observations, and $\bar{\tilde{y}}_{i,.}$ indicates the time-series average of the head’s income. In this cross-sectional model, I include a set of family composition variables to control for unobservable preferences because the household-fixed effect cannot be modeled as an error component. If retailers allowed households to use credit purchases regardless of the variability of earnings, the estimate of $\varkappa_{1}$ becomes close to zero. By contrast, if retailers did not allow households with high-income volatility to trade on credit, the estimate should be positive and reasonably large. Online Appendix C.5 summarizes the results. Shortly, the estimated coefficient ($\hat{\varkappa}_{1}$) is close to zero and statistically insignificant in different specifications. This result remains unchanged when alternative thresholds for the dependent indicator variable are used to include the households with less than the first quartile of the average credit purchases. This suggests that retailers may be generous in considering the volatility of earnings when providing credit to factory workers, which is consistent with the retailers’ incentive to secure their local customers.
@@
This, in turn, implies that retailers may factor risk into retail prices or by charging interest on their loans. Unfortunately, there are no systematic statistics on the interest charged on credit purchases. It may be possible to reverse calculate the risk premiums by differencing payments and purchases. This reverse calculation technique requires information on both the price and type of retailers. However, budget books usually document price-by-item information but not price-by-retailer information. This is problematic because the same items were often sold in different types of retailers: for example, eggs were frequently sold in greengrocers (yaoya) but sometimes also in general stores (aramonoya). Fortunately, I found a specific month in a budget book that has documented the price-by-retailer information. [^71] The OISR, Archives of the Tsukishima Survey (THBS #4, February 1919). According to this budget book, the reimbursements were made semi-monthly, as seen in Figure 6b. They bought various goods in greengrocers, fish stores, and general stores between February 1 and February 14, 1919, and paid exactly the same amounts on the 15th. [^72] The repayment period in credit purchases among retailers and wholesalers (footnote 76) was often set at approximately one month (Tokyo City Office 1932, pp. 157-159). Therefore, customers likely made repayments at semi-monthly or monthly intervals. In the latter half of February, the household paid the same amounts to greengrocers and general stores on the 28th, but paid less to the fish stores. As the budget book for March does not have information on the type of retailers, it is difficult to determine exactly how the reimbursement shortfall was managed. However, this case suggests that the risk premiums may have been included in the retail prices but not in the form of interest.

### v14

@@ 4.3 Response to Adverse Health Shocks
The identification in the analyses above relies on both adverse and favorable idiosyncratic shocks. To focus on the mechanisms behind consumption smoothing in response to adverse shocks, I use short-term income losses due to the temporary illnesses of household heads for the identification in this section.
@@
Notes: This figure shows time-series plots of the monthly death counts from influenza (ryūkōsei kanbō) in Tokyo prefecture from January 1918 to December 1920. The death statistics are collected through the comprehensive registration system. Source: Created by the author using the 1918, 1919, and 1920 editions of the Statistics of Causes of Deaths of the Empire of Japan (Statistics Bureau of the Cabinet 1921; 1922; 1923).
@@
I begin my analysis by testing whether the head’s illness affected earnings. Figure 4 shows the monthly death counts from influenza (ryūkōsei kanbō) in Tokyo between 1918 and 1920. Early 1919 saw a moderate spike, partially captured in the THBS period. Indeed, a few THBS budget books indicate that family members contracted influenza during this time. According to the Kazoku idō hyō (Section 3.4), five households experienced illness during the sample period, with illness affecting the heads concentrated between February and May 1919, when influenza exposure was most prevalent. The average duration of illness for the heads was $5.8$ days (Std. Dev. =3.6=3.6), consistent with Type A influenza, which causes fever lasting less than six days (Uyeki et al. 2022). Therefore, influenza likely caused the illnesses during this period.
@@
Earnings in a semi-month are based on work performed during the previous semi-month. Table 6 presents the regression results of the head’s earnings on the number of illness days in the previous semi-monthly cell. While instantaneous health shocks have little impact on earnings (Column 1), lagged health shocks have a statistically significant effect (Column 2). This result remains unchanged when both variables are included in the same regression (Column 3), supporting the evidence that health shocks in the previous cell reduce the head’s earnings. The estimate in Column 2 suggests that a one day increase in illness reduces the head’s semi-monthly income by approximately $2.4$ yen. Notably, this is close to the average daily wage of skilled factory workers, reflecting salary reductions due to absence (Online Appendix B.6). This aligns with anecdotal evidence that the head’s illness directly caused income loss due to absence (Miyoshi 1989). Finally, Column 4 adds the health shocks of other family members to the specification in Column 3. The estimated coefficients on these placebo variables are statistically insignificant, supporting the robustness of the results in Column 3.
@@
Notes: This table presents results from regressions of the head’s earnings on health shock variables using the adjusted semi-monthly panel dataset. Households with only two semi-month observations are excluded from all regressions because lagged health shock variables are included. All regressions include the family size control, household fixed effects, and time fixed effects. “Health shocks on the head”: the number of days the head was ill. “Health shocks on others”: the number of days other family members (excluding the head) were ill. “[level]”: the variable measured in the same semi-month cell. “[lag]”: the variable from the previous semi-month cell. Although the correlation between the level and lag of health shocks for other family members is high (Pearson’s coefficient =0.99=0.99), including either variable does not change the finding that health shocks for other family members have negligible effects.
I then characterize equation 1 as a structural equation by using the first-stage reduced form equation as follows:
@@
where $h_{i,t-1}$ is the number of days the head was ill in the previous semi-month cell, $\varpi_{i}$ is the household fixed effect, $\tau_{t}$ is the semi-month fixed effect, and $e_{i,t}$ is a random error term. For the analysis using equation 2, the log-transformed disposable income in equation 3 ($y_{i,t}$) is replaced with the head’s earnings ($\tilde{y}_{i,t}$). The health shock variable is plausibly exogenous because the illnesses resulted from influenza epidemics. [^50] As noted in Section 2.3, factories did not provide sick leave compensation because the Factory Act did not specify benefits for workers’ illness (Study Group of the Factory Act 1916). Even at the largest factory in Tsukishima, the Ishikawajima Shipyard, the company did not record any compensation for influenza epidemics (Arai 1930). In addition, the health shock variable is statistically correlated with the head’s earnings (Table 6). Together, this evidence supports the exogeneity of the instrument and satisfies the relevance condition in this just-identified system.
@@
Notes: Column 1 reports results of equation 1, replacing log-transformed disposable income with the head’s illness (number of days ill in the previous semi-month). Dependent variables are the 11 measures of log-transformed consumption. All regressions include the family size control, household fixed effects, and time fixed effects. The coefficient on the health shock variable is shown in the “Coef.” sub-column. Column 2 treats log-transformed disposable income as endogenous, using the head’s illness as an instrument in the first-stage equation (3). The estimated coefficient on disposable income appears in the “Coef.” sub-column, and first-stage $F$-statistics for the instrument are listed in the final column.
@@
I first consider a reduced-form specification using the head’s illness as the exposure variable in equation 1, following Cochrane (1991). Column 1 in Table 7 presents the results, providing evidence that adverse health shocks to the heads reduced consumption levels. Next, I explicitly model the structure underlying the relationship between health shocks and earnings using the reduced-form equation in (3). If the negative impacts of health shocks on consumption derive from income losses, the results from the instrumental variable approach should correspond to those in Column 1. The estimates are listed in Column 2 of Table 7. As shown, these results align with the reduced-form findings in Column 1. Importantly, the estimates are systematically larger than those reported in Table 4, with total and food consumption estimates of $0.543$ and $0.355$, respectively. While this partly reflects bias reduction under instrumental variable estimation, it also reflects the system’s design to reduce bias by specifying the channel from adverse (rather than favorable) income shocks. For example, allowances for children can be cut during unfavorable periods but may not increase during favorable ones. Similarly, furniture expenses are more likely to be reduced during shocks, but households purchase furniture infrequently, providing few observations for rare health shocks. [^51] In other words, furniture may only have been purchased when households had extra income. This pattern may partly explain the smaller first-stage $F$-statistics for the furniture subcategory. [^52] Note that the logarithmic transformation retains cells with furniture expenses. Thus, the reduced-form equation captures the relationship between health shocks and earnings in cells with any furniture expenses. While the first-stage $F$-statistic is $35.7$ in the linear-linear model, the second-stage result is robust to the functional form assumption. This may reflect the asymptotic advantage of panel data models under nearly weak instruments (Staiger and Stock 1997; Cai et al. 2012; Baltagi et al. 2012). The same argument applies to the housing category. A similar interpretation applies to the transportation category, as family travel is infrequent and unlikely to coincide with illness events. By contrast, entertainment expenses are still likely to be cut in case of illness, reflecting reductions in tobacco use, theater visits, and allowances for heads. This may explain the differences from the baseline results in Table 4.
@@
Although the estimates for clothes and gifts now exceed one, the null hypothesis of equality to the population mean of one is not statistically rejected for either category. While the large estimates may partly result from the instrument based on rare health shock events, the findings are broadly consistent with the baseline results—both categories remain sensitive to shocks, similar to other dispensable categories.
@@
Adverse Health as the Idiosyncratic Shock
@@
Head’s illness Obs. Head’s earnings Obs.
@@
Panel A: Net savings, insurance, borrowing, and gifts (income - expenditure, yen)
@@
Other members’ earnings 0.001 [0.081] 254 -0.001 [0.034] 254
@@
Notes: Column 1 reports the results of equation 2, where the head’s earnings are replaced by the head’s illness (i.e., the number of days ill in the previous semi-month cell). Family size, household fixed effects, and time fixed effects are included in all regressions. The estimated coefficients on the head’s illness are listed in the sub-column labeled “Coef.” Column 2 reports the results of equation 2, treating the head’s earnings as the endogenous variable. The head’s illness serves as an instrumental variable in the first-stage reduced-form equation. The first-stage results match those in Column 2of Table 6, with a first-stage $F$-statistic of $13.6$. The estimated coefficients on the head’s earnings are listed in the sub-column labeled “Coef.” Panel A: Each net income variable is defined as income minus expenditure. “Net savings”: withdrawals minus deposits. “Net insurance”: insurance receipts minus premiums paid. “Net borrowing”: borrowing minus debt repayments. This category considers net income from both pawnshops and other lending institutions. “Net credit purchase”: total credit purchases minus repaid credit. “Net gifts”: total received gifts, including pecuniary and non-pecuniary, minus any gift payments. Panel B: “Other members’ earnings”: the total income earned by all household members except the head. “Sales of miscellaneous assets”: include sales of daily items such as newspapers and empty bottles. All dependent variables are in yen.
@@
Table 8 presents the analysis of the risk-coping strategies for mitigating adverse health shocks. Column 1 presents the reduced-form results. The estimated coefficients on the net income variables are positive in most cases, indicating households’ responses to the shocks. The estimate for the credit purchase category is weakly statistically significant and has the largest magnitude. This suggests that a one standard deviation increase in the duration of the head’s flu exposure increases net credit purchases by $4.6$ yen ($1.29\times 3.6~\text{days}$), which represents approximately 1616% of their average semi-monthly earnings of the head (Table 3). Column 2 reports the estimates from the instrumental variable approach, treating the head’s earnings as endogenous. The estimate for the credit purchase category is approximately four times larger than that reported in Column 1 of Table 5. It implies that a one yen decrease in the head’s earnings increases net income from credit purchases by $0.55$ yen. This magnitude aligns with the direct effect of health shocks on net credit purchases estimated in Column 1 of Table 8.535353Given their average daily wage of $2.56$ (Online Appendix B.6), $2.56\times 0.55=1.4$ yen, which is close to $1.3$ yen in Column 1. The estimates for the savings and gifts categories remain negative but are now statistically insignificant.
Overall, these results suggest that the primary response to adverse idiosyncratic income shocks among THBS households was to use credit purchases. Compensation via credit purchases accounted for roughly half of their temporary income losses, helping to smooth consumption in the short term.
5 Credit System in the Local Economy
@@
The results highlight the importance of the relationships between workers and local retailers. To the best of my knowledge, however, little historical research exists on credit purchases in prewar Japan, and the details of this credit institution remain unclear. As with state-contingent loans in the rural credit market (Udry 1994), information on credit purchases is personal and poorly documented. Accordingly, I analyze the institution using the THBS dataset along with a few available official documents.
@@ 5.1 Availability and Premiums: Evidence from the THBS Dataset
Notes: Figures 5a and 5c plot average daily and semi-monthly credit purchases and payments, respectively. The blue and red dashed lines indicate the average credit purchases and payments, respectively. Figures 5b and 5d plot credit purchases minus payments. The red solid line marks zero. Daily data cover January 12–November 11, 1919. Figures 5c and 5d show the semi-monthly series derived from the same daily data. All figures use the 29 THBS households that made credit purchases. Source: Created by the author using the THBS sample.
I begin this descriptive analysis by examining credit purchases in the THBS dataset. Most THBS households ($29/33$) made credit purchases during the sample period, confirming that this practice was common among factory workers. Figure 5a illustrates the average daily credit purchases and payments. Payments generally peak semi-monthly, reflecting the practice of settling credit purchases on or shortly after paydays. The purchase series shows a similar pattern but with smaller peaks and greater variance. This indicates that households first make purchases on or soon after paydays and then consume perishable goods such as fish, fruit, and vegetables. Figure 5b, which illustrates net purchases, confirms this behavior: households used credit daily and repaid it near paydays. Figure 5c shows that payments slightly exceed purchases, implying near-full loan repayment for most households. Although the difference between payment and purchase becomes more noticeable after August 1919, it is not statistically significant in all semi-months (Figure 5d). [^54] Strictly, payments can include accrued liabilities from prior periods. Therefore, average payments reasonably exceed average purchases in most semi-month cells. Note that the wider variances in the later sample period are owing to the smaller number of available households (Online Appendix Figure B.4).
Next, I analyze whether earnings volatility restricts credit purchases. Because household credit is unobservable, I use earnings volatility as a proxy for credit risk in the following empirical specification:
@@
where $I(\cdot)$ is an indicator function, $\overline{CP}$ is the (unit-)average credit purchases, $CV$ is the coefficient of variation for the head’s earnings, $\bar{\mathbf{x}}$ is a vector of (unit-)average family composition variables, and $v$ is a random error term. All variables are calculated using the semi-monthly series. [^55] For example, the coefficient of variation for household $i$ is defined as $CV_{i}=\{\sum_{t}^{T_{i}}(\tilde{y}_{i,t}-\bar{\tilde{y}}_{i,.})^{2}/(T_{i}-1)\}^{\frac{1}{2}}/\bar{\tilde{y}}_{i,.}$, where $T_{i}$ indicates the number of time-series observations, and $\bar{\tilde{y}}_{i,.}$ is the time-series average of the head’s income. In this cross-sectional model, I include family composition variables to control for unobserved preferences, since household fixed effects cannot be included as an error component. If retailers allow households to use credit regardless of earnings variability, the estimate of $\varkappa_{1}$ should be close to zero. By contrast, if households with high earnings volatility are restricted from trading on credit, the estimate should be positive and sizable. Online Appendix C.7 summarizes the results. In short, the estimated coefficient ($\hat{\varkappa}_{1}$) is close to zero and statistically insignificant. This finding remains unchanged when alternative thresholds for the dependent indicator variable are applied, including households below the first quartile of average credit purchases. This suggests that retailers were generous in considering earnings volatility when extending credit to factory workers, consistent with their incentive to retain local customers (Section 5.2).
@@
This, in turn, suggests that retailers may factor risk into retail prices or by charging interest on loans. Unfortunately, no systematic statistics exist on interest charged for credit purchases. One might attempt to reverse-calculate risk premiums by differencing payments and purchases. However, this approach requires information on both the price and type of retailer. Budget books usually record prices by item, but not by retailer, which is problematic because the same items are often sold by different types of retailers–for example, eggs are commonly sold in greengrocers (yaoya) but occasionally also in general stores (aramonoya). Fortunately, a specific month in a budget book provides price-by-retailer information. [^56] The OISR, Archives of the Tsukishima Survey (THBS #4, February 1919). According to this record, reimbursements were made semi-monthly, as shown in Figure 5b. Between February 1 and February 14, 1919, the household purchased various goods from greengrocers, fish stores, and general stores, and paid exactly the same amounts on the 15th. [^57] The repayment period for credit purchases among retailers and wholesalers (footnote 61) was often set at approximately one month (Tokyo City Office 1932, pp. 157-159). Therefore, customers likely made repayments at semi-monthly or monthly intervals. In the latter half of February, the household paid the same amounts to greengrocers and general stores on the 28th, but less to fish stores. Because the March budget book lacks information on retailer type, determining exactly how the reimbursement shortfall was handled is difficult. Nevertheless, this case suggests that risk premiums were incorporated into retail prices rather than charged as explicit interest.